The question is - to prepare 5 milimeters of a 10 millimolar solution of glucose from a 1 molar stock solution how many micro litres of stock need to be added to how many millilitres of water


1
Expert's answer
2021-03-18T05:57:23-0400

5milimetres = (10 millimolar)(1) = (10)(5)

= 2microlitres

= 2000 microlitres

10 -2 = 8 milimetres of water

= 0.1×8 = 0.8 milliliters of water